one canned juice drink is 25% orange juice; another is 5%orange juice. How many liters of each should be mixed together in order to get 20L that is 19% orange juice
Answer by mananth(16949)

25% orange juice; let it be x liters
another is 5%orange juice. this will be 20-x liters
20L that is 19% orange juice
...
0.25x+0.05(20-x) = 0.19*20
0.25x+1-0.05x=1.8
0.20x=1.8-1
0.20x=0.8
x=4 liters of 25% juice
16 liters of 5% juice

m.ananth@hotmail.ca